OpenStudy (anonymous):

Which point lies on the line 3x+7y=2 ?

OpenStudy (anonymous):

Well, since there a line is a function, for every X there will be a Y. A point serious is defined by an (X,Y) patter. Change it to linear format: 3x+6y=2 6y= - 3x+2 y = -(3/6)x+2/6 y= -.5x+ 1/3 So pick an x and put that number in and you get a y For exmaple: X = 0 y = -.5x+ 1/3 y= .5(0) + 1/3 y= 0 + 1/3 y=1/3 So the solution is: (0, 1/3)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No problem , you can thank me by medaling me c; lol